OpenStudy (anonymous):

6x+9/15x^2 divided by 16x-12/10x^4 options: a) x^2(2x+3)/2(4x-3) b)2x+3/2(4x-3) c)2(4x-3)/x^2(2x+3) d) 8x^2(4x-3)/9(2x+3)

OpenStudy (turingtest):

Please try to explain your answer a little @CoCoTsoi

OpenStudy (anonymous):

6x+9/15x^2 divided by 16x-12/10x^4 =(6x+9)(10x^4)/((15x^2) (16x-12)) =(2x+3)(x^2) / 2(4x-3) =x^2(2x+3)/2(4x-3)

OpenStudy (anonymous):

It was correct- tysm!!